# Hyper-Compression: Model Compression via Hyperfunction

### v3

@@ Abstract
The rapid growth of large models’ size has far outpaced that of computing resources. To bridge this gap, encouraged by the parsimonious relationship between genotype and phenotype in the brain’s growth and development, we propose the so-called hyper-compression that turns the model compression into the issue of parameter representation via a hyperfunction. Specifically, it is known that the trajectory of some low-dimensional dynamic systems can fill the high-dimensional space eventually. Thus, hyper-compression, using these dynamic systems as the hyperfunctions, represents the parameters of the target network by their corresponding composition number or trajectory length. This suggests a novel mechanism for model compression, substantially different from the existing pruning, quantization, distillation, and decomposition. Along this direction, we methodologically identify a suitable dynamic system with the irrational winding as the hyperfunction and theoretically derive its associated error bound. Next, guided by our theoretical insights, we propose several engineering twists to make the hyper-compression pragmatic and effective. Lastly, systematic and comprehensive experiments confirm that hyper-compression enjoys the following PNAS merits: 1) Preferable compression ratio; 2) No post-hoc retraining; 3) Affordable inference time; and 4) Short compression time. It compresses LLaMA2-7B in an hour and achieves close-to-int4-quantization performance, without retraining and with a performance drop of less than 1%. We have open-sourced our code in https://github.com/Juntongkuki/Hyper-Compression.git for free download and evaluation.
keywords:
@@ 1 Introduction
Recently, due to the pursuit of the scaling law, the escalating demand for computational resources by large models has presented formidable challenges for their deployment in resource-constrained environments Ding et al. (2023). For example, a GPT-3 model has 175 billion parameters with a size of approximately 700 GB, while one of the most advanced GPUs (NVIDIA H100) has a memory capacity of only up to 80 GB. Serving large models well has become a strong technological imperative. To this end, currently, one prevalent way is to develop effective model compression approaches that crop the size of models while maintaining the performance.
@@
Model compression Zhu et al. (2023) predominantly revolves around four different classes of algorithms: pruning, quantization, knowledge distillation, and low-rank decomposition. All these algorithms were already proposed years or even decades ago. Though these techniques play a critical role in model compression, they face intrinsic challenges in the era of large models that are hard to overcome. First, the compression efficacy of some methods such as pruning and quantization is challenging to scale. A plethora of studies showed that pruning is submitted by the suboptimal compression rate Zhu et al. (2023), usually 24×2-4$\times$2 - 4 ×. As for quantization, even the best it can do (1-bit quantization) remains unexciting because the compression ratio is bounded by a constant, which is handicapped to address the enlarged gap between models and the hardware resource in the era of large models. Moreover, when compressing a model at a large rate, the model’s output is usually severely distorted. Thus, retraining and recalibration are mandated to recover the model’s performance, which adds extra costs for curating data and computational resources. This not only is unfriendly for the industry that often favors agile deployment but also may introduce bias in the model. Observing these fundamental limits, we ask can we have a more promising technology roadmap, instead of keeping modifying the existing methods? The answer is affirmative. Here, we introduce hyper-compression, a novel and general-purpose approach that redefines model compression as a problem of parsimonious parameter representation. This concept stems from hyper-networks that are smaller networks but can generate weights for a significantly larger target network Stanley et al. (2009); Chauhan et al. (2024), in analogy to the famous genomic bottleneck that a comparatively small genome can control the growth and development of a complicated brain Shuvaev et al. (2024). We generalize the concept of hypernets into a ‘hyperfunction’ (HF) and hypothesize that weights of a large network can be encoded by a function with few parameters. Mathematically, we use a parametric function $w_{n}=h(\theta;n)$ to encode the relationship between locations and weights of the target network, where $w_{n}$ is the $n$-th parameter of the target network, and $\theta$ collects the parameters of $h$. Should the memory footprint of $\theta$ be significantly less than that of $\{w_{n}\}_{n=1}^{N}$, $\theta$ can be stored on devices as a kind of compression for network parameters. During inference, $\theta$ is freely used to layer-by-layer recover weights of the original network via $h$.
@@
Hereafter, we refer to model compression via the hyperfunction as the hyper-compression. Along this direction, we reasonably underscore that finding a suitable hyperfunction is an open-ended question. However, unexpectedly, we propose to leverage the density of trajectory in ergodic theory Cornfeld et al. (2012) to construct the hyper-function $h$. One of the most important findings in ergodic theory is that when the transformation $T$ is ergodic, a low-dimensional dynamic system can eventually visit a high-dimensional space given sufficient time. Hereafter, we refer to this property as the trajectory density. Informally, there exists $\boldsymbol{x}_{0}$ such that ${\forall\,}\boldsymbol{x}$ and $\epsilon>0$, ${\exists\,}k$ such that
$\|\boldsymbol{x}-\underbrace{(T\circ T\circ\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu}\circ T)}_{k}(\boldsymbol{x}_{0})\|=\|\boldsymbol{x}-T^{k}(\boldsymbol{x}_{0})\|<\epsilon.$ (1)
Thus, the hyperfunction $h=T^{k}(\cdot)$ and we can “compress” a high-dimensional vector $\boldsymbol{x}$ into a number $k$, which should induce a good compression rate. Next, as the initial step, we use the irrational winding Katok et al. (1995) as the suitable dynamic system to validate the feasibility of the proposed framework for model compression.
Then, we theoretically derive the error bound of the hyper-compression based on the irrational winding, thereby characterizing the key ingredients that affect the error. Next, guided by theoretical insight, we propose several pragmatic engineering twists such as scaling to cope with outliers appropriately, adjusting irrational direction to diminish the error, and applying the KD tree Zhou et al. (2008) for fast compression. These twists lead to an effective and efficient model compression algorithm. Specifically, it offers four distinct characters, succinctly summarized as PNAS: 1) Preferable compression ratio; 2) No post-hoc retraining&recalibration; 3) Affordable inference time; and 4) Short compression time. Our work can facilitate the harmony between the scaling law and the stagnation of hardware upgradation in terms of saving both computation and data. More favorably, our method is seamlessly compatible with other compression methods such as pruning, distillation, and low-rank decomposition to further amplify the compression efficacy. Systematic and comprehensive experiments confirm the efficacy of hyper-compression. In summary, our contributions are threefold:
@@
iii) Systematic and comprehensive experiments on large models such as the LLaMA et al. (2023) series and small models such as UNet Ronneberger et al. (2015) and Mobile-Net Howard et al. (2019) confirm the competitiveness of the proposed hyper-compression, which is friendly to agile deployment in both academia and industry.
@@ 2.1 Model Compression
Model compression Zhu et al. (2023) plays a crucial role in the widespread deployment of deep learning models in a myriad of different settings. There are mainly four different classes of techniques including pruning, quantization, low-rank decomposition, and knowledge distillation. As Table 1 shows, we highlight that hyper-compression is a fundamentally different method that performs the low-dimensional transformation for the target data. Such an essential difference can open a lot of doors for research and translation opportunities. We summarize model compression techniques below, accompanied by the recent advances in compressing large language models. Due to the limit of pages, we can only cover a few representative works. Moreover, because our proposed hyper-compression here focuses on post-training compression, and knowledge distillation is to a large extent based on training, we do not review articles on knowledge distillation here.
Table 1: Hyper-compression compresses a model based on an essentially different mechanism, which performs the low-dimensional transformation for the target data.
@@
Unstructured pruning targets individual parameters, without the need of considering the internal structures of a model. SparseGPT Frantar and Alistarh (2023) turned the pruning problem into a set of extremely large-scale instances of sparse regression, thereby avoiding the inversion of each matrix. Thus, SparseGPT could compress models of 10-100 billion parameters in just a few hours. Wanda Sun et al. (2023) simultaneously considers weights and activations in pruning. This technique was motivated by an observation in Dettmers et al. (2022) that a small subset of hidden features are substantially large in magnitude. Therefore, they augmented the standard weight magnitude with the input activations to evaluate the importance of weight.
@@
Structured pruning usually removes the entire neurons, filters, and blocks of a network, thereby leading to realistic compression and acceleration. Unlinke the conventional methods, EBert Liu et al. (2021) incorporated a predictor to dynamically pinpoint and remove unimportant heads in multi-head self-attention layers and unimportant structured computations in fully-connected networks, respectively, when inferring each batch of samples. LLM-shearing Xia et al. (2023) found that the pruned model has an imbalanced performance across different domains and tasks, and proposed to dynamically load data batch from each domain in proportion to its rate of performance reduction in that domain. K-Prune Park et al. (2023), a retraining-free structured pruning method, used the knowledge loss to measure the importance of heads in transformers, where the knowledge loss is defined as the faithfulness to the soft labels of the original model. FLAP An et al. (2024) found that certain channels of hidden state features exhibit structured sample stability, and designed a structured pruning metric to identify whether the output feature map is easy to recover when a column of weight matrix is eliminated. Then, FLAP adds an additional bias term to recover the output feature.
@@
Quantization reduces the precision of weights and activations in a neural network by turning the high bit-width numbers into lower bit-width ones. Quantization can directly diminish memory usage and accelerate computations by using fixed-point arithmetic. It enables faster inference on hardware with limited computational resources. Quantization entails quantization-aware training (QAT Liu et al. (2023)) and post-training quantization (PTQ Huang et al. (2024)). The former emphasizes the combination of quantization and training. The latter quantizes parameters when the network training is completed, which does not require modifications to the network structure but may induce the extra cost of retraining. Therefore, here we mainly discuss PTQ, which is divided into weight quantization and weight&activation quantization.
@@
LLM.int8() Dettmers et al. (2022) initially employs vector-wise quantization alongside distinct normalization constants for every inner product within the matrix multiplication to quantize most of the features. Moreover, LLM.int8() denotes outlier feature dimensions by a 16-bit matrix multiplication, while over 99.9% of values are still multiplied in 8-bit. OPQ Frantar and Alistarh (2022) is the layer-wise compression method, which performs quantization layer-by-layer and minimizes the error between the pre-activation by the original full-precision and the quantized weight matrix in each layer. SmoothQuant Xiao et al. (2023) is a training-free, accuracy-preserving, and general-purpose post-training quantization solution. It smoothens the activation outliers by a mathematical scaling transformation before the normal quantization:
@@
where $X$ is the activation and $W$ is the weight matrix. AWQ Lin et al. (2024) highlights that not all weights are equally crucial for large language model (LLM) performance. Only 0.1%-1% of weights are significant; bypassing quantization for these salient weights can greatly reduce quantization loss. To identify these salient channels, they examined activation distributions rather than weight distributions. To avoid inefficient mixed-precision implementations, they analyzed weight quantization errors and determined that scaling up salient channels can decrease their relative quantization error. Atom Zhao et al. (2024) also designs a customized CUDA kernel that utilizes low-bit tensor cores, besides the common quantization operations such as mixed precision and grouped quantization. Atom improves end-to-end throughput (token/s) by up to $7.73$ times compared to the FP16.
@@
Low-Rank Adaptation (LoRA, Hu et al. (2022)) posits that the adjustments in weights during model adaptation exhibit a low ”intrinsic rank.” This approach enables the training of certain dense layers in a neural network indirectly by optimizing the gradients into a low-rank decomposition matrix in those layers during adaptation, while keeping the pre-trained weights unchanged. TensorGPT Xu et al. (2023) introduced the approach of tensorizing and decomposing each token embedding, rather than treating the entire embedding matrix as a single entity. It was the first method to apply the tensor train decomposition for model compression, achieving a reduction in the number of parameters by a factor of 2.31.
@@ 2.2 Implicit Neural Representation
Implicit Neural Representation (INR) Park et al. (2019) has emerged as a transformative approach in computer graphics, computer vision, and machine learning. This technique leverages neural networks to represent complex shapes and images without explicitly storing the geometry or pixel data. Early work in this area focused on using neural networks to learn mappings from coordinates to values. For instance, the pioneering work by Park et al. (2019) introduced the concept of using multi-layer perceptrons (MLPs) to represent 3D shapes as continuous functions, allowing for high-resolution representations without traditional mesh-based methods. Recently, the implicit neural representation is increasingly used for data compression, such as COIN Dupont et al. (2021) and NERV++ Ghorbel et al. (2024).
@@ 2.3 Hypernet
The concept of hypernetworks was first articulated by Ha et al. (2017), who proposed a framework where a neural network (the hypernetwork) generates the weights of another network. This meta-learning approach allows for dynamic adaptation, enabling a single hypernetwork to cater to multiple tasks by producing tailored weights. The hypernetwork architecture is particularly advantageous for scenarios where training multiple models is computationally expensive or infeasible. Hypernet as a meta-learning method has been successfully applied in a plethora of fields including few-shot learning Rusu et al. (2018) and domain adaptation Volk et al. (2022), to name a few.
@@ 3.1 Hyper-Compression
A small human genome can remarkably encode the development of a human brain to the scale of billions of neurons and trillions of synaptic connections Stanley et al. (2009); Shuvaev et al. (2024). This observation suggests the existence of an implicit mapping from genotype to phenotype, capable of expanding a limited number of genetic instructions into a vast array of biological substances. Inspired by this efficient genetic encoding, the hypernet Chauhan et al. (2024) uses a small network (genotype, called hypernet) to control the design of the target network (phenotype) including architectures and weight distributions.

### v4

@@ Abstract
The rapid growth of large models’ size has far outpaced that of computing resources. To bridge this gap, encouraged by the parsimonious relationship between genotype and phenotype in the brain’s growth and development, we propose the so-called Hyper-Compression that turns the model compression into the issue of parameter representation via a hyperfunction. Specifically, it is known that the trajectory of some low-dimensional dynamic systems can fill the high-dimensional space eventually. Thus, Hyper-Compression, using these dynamic systems as the hyperfunctions, represents the parameters of the target network by their corresponding composition number or trajectory length. This suggests a novel mechanism for model compression, substantially different from the existing pruning, quantization, distillation, and decomposition. Along this direction, we methodologically identify a suitable dynamic system with the irrational winding as the hyperfunction and theoretically derive its associated error bound. Next, guided by our theoretical insights, we propose several engineering twists to make the Hyper-Compression pragmatic and effective. Lastly, systematic and comprehensive experiments on NLP models such as LLaMA and Qwen series and vision models confirm that Hyper-Compression enjoys the following PNAS merits: 1) Preferable compression ratio; 2) No post-hoc retraining; 3) Affordable inference time; and 4) Short compression time. It compresses LLaMA2-7B in an hour and achieves close-to-int4-quantization performance, without retraining and with a performance drop of less than 1%. We have open-sourced our code in https://github.com/Juntongkuki/Hyper-Compression.git for free download and evaluation.
Index Terms:
@@ 1 Introduction
Recently, due to the pursuit of the scaling law, the escalating demand for computational resources by large models has presented formidable challenges for their deployment in resource-constrained environments [15]. For example, a GPT-3 model has 175 billion parameters with a size of approximately 700 GB, while one of the most advanced GPUs (NVIDIA H100) has a memory capacity of only up to 80 GB. Serving large models well has become a strong technological imperative. To this end, currently, one prevalent way is to develop effective model compression approaches that crop the size of models while maintaining the performance.
@@
Model compression [62] predominantly revolves around four different classes of algorithms: pruning, quantization, knowledge distillation, and low-rank decomposition. All these algorithms were already proposed years or even decades ago. Though these techniques play a critical role in model compression, they face intrinsic challenges in the era of large models that are hard to overcome. First, the compression efficacy of some methods such as pruning and quantization is challenging to scale. A plethora of studies showed that pruning is submitted by the suboptimal compression rate [62], usually $2-4\times$. As for quantization, even the best it can do (1-bit quantization) remains unexciting because the compression ratio is bounded by a constant, which is handicapped to address the enlarged gap between models and the hardware resource in the era of large models. Moreover, when compressing a model at a large rate, the model’s output is usually severely distorted. Thus, retraining and recalibration are mandated to recover the model’s performance, which adds extra costs for curating data and computational resources. This not only is unfriendly for the industry that often favors agile deployment but also may introduce bias in the model. Observing these fundamental limits, we ask can we have a more promising technology roadmap, instead of keeping modifying the existing methods? The answer is affirmative. Here, we introduce Hyper-Compression, a novel and general-purpose approach that redefines model compression as a problem of parsimonious parameter representation. This concept stems from hyper-networks that are smaller networks but can generate weights for a significantly larger target network [46, 9], in analogy to the famous genomic bottleneck that a comparatively small genome can control the growth and development of a complicated brain [45]. We generalize the concept of hypernets into a ‘hyperfunction’ (HF) and hypothesize that weights of a large network can be encoded by a function with few parameters. Mathematically, we use a parametric function $w_{n}=h(\theta;n)$ to encode the relationship between locations and weights of the target network, where $w_{n}$ is the $n$-th parameter of the target network, and $\theta$ collects the parameters of $h$. Should the memory footprint of $\theta$ be significantly less than that of $\{w_{n}\}_{n=1}^{N}$, $\theta$ can be stored on devices as a kind of compression for network parameters. During inference, $\theta$ is freely used to layer-by-layer recover weights of the original network via $h$.
@@
Hereafter, we refer to model compression via the hyperfunction as the Hyper-Compression. Along this direction, we reasonably underscore that finding a suitable hyperfunction is an open-ended question. However, unexpectedly, we propose to leverage the density of trajectory in ergodic theory [12] to construct the hyperfunction $h$. One of the most important findings in ergodic theory is that when the transformation $T$ is ergodic, a low-dimensional dynamic system can eventually visit a high-dimensional space given sufficient time. Hereafter, we refer to this property as the trajectory density. Informally, there exists $\bm{x}_{0}$ such that $\forall\bm{x}$ and $\epsilon>0$, $\exists k$ such that
$\|\bm{x}-\underbrace{(T\circ T\circ\cdots\circ T)}_{k}(\bm{x}_{0})\|=\|\bm{x}-T^{k}(\bm{x}_{0})\|<\epsilon.$ (1)
Thus, the hyperfunction $h=T^{k}(\cdot)$ and we can “compress” a high-dimensional vector $\bm{x}$ into a number $k$, which should induce a good compression rate. Next, as the initial step, we use the irrational winding [29] as the suitable dynamic system to validate the feasibility of the proposed framework for model compression.
Then, we theoretically derive the error bound of the Hyper-Compression based on the irrational winding, thereby characterizing the key ingredients that affect the error. Next, guided by theoretical insight, we propose several pragmatic engineering twists such as scaling to cope with outliers appropriately, adjusting irrational direction to diminish the error, and applying the KD-Tree [61] for fast compression. These twists lead to an effective and efficient model compression algorithm. Specifically, it offers four distinct characters, succinctly summarized as PNAS: 1) Preferable compression ratio; 2) No post-hoc retraining&recalibration; 3) Affordable inference time; and 4) Short compression time. Our work can facilitate the harmony between the scaling law and the stagnation of hardware upgradation in terms of saving both computation and data. More favorably, our method is seamlessly compatible with other compression methods such as pruning, distillation, and low-rank decomposition to further amplify the compression efficacy. Systematic and comprehensive experiments confirm the efficacy of Hyper-Compression. In summary, our contributions are threefold:
@@
iii) Systematic and comprehensive experiments on large models such as the LLaMA [17] series, Qwen [1] series and small models such as ResNet [25] series, UNet [42] and Mobile-Net [26] confirm the competitiveness of the proposed Hyper-Compression, which is friendly to agile deployment in both academia and industry.
@@ 2.1 Model Compression
Model compression [62] plays a crucial role in the widespread deployment of deep learning models in a myriad of different settings. There are mainly four different classes of techniques including pruning, quantization, low-rank decomposition, and knowledge distillation. As Table I shows, we highlight that Hyper-Compression is a fundamentally different method that performs the low-dimensional transformation for the target parameters. Such an essential difference can open a lot of doors for research and translation opportunities. We summarize model compression techniques below, accompanied by the recent advances in compressing large language models. Due to the limit of pages, we can only cover a few representative works. Moreover, because our proposed Hyper-Compression here focuses on post-training compression, and knowledge distillation is to a large extent based on training, we do not review articles on knowledge distillation here.
TABLE I: Hyper-Compression compresses a model based on an essentially different mechanism, which performs the low-dimensional transformation for the target data.
@@
Unstructured pruning targets individual parameters, without the need of considering the internal structures of a model. SparseGPT [19] turned the pruning problem into a set of extremely large-scale instances of sparse regression, thereby avoiding the inversion of each matrix. Thus, SparseGPT could compress models of 10-100 billion parameters in just a few hours. Wanda [47] simultaneously considers weights and activations in pruning. This technique was motivated by an observation in [14] that a small subset of hidden features are substantially large in magnitude. Therefore, they augmented the standard weight magnitude with the input activations to evaluate the importance of weight.
@@
Structured pruning usually removes the entire neurons, filters, and blocks of a network, thereby leading to realistic compression and acceleration. Unlinke the conventional methods, EBert [36] incorporated a predictor to dynamically pinpoint and remove unimportant heads in multi-head self-attention layers and unimportant structured computations in fully-connected networks, respectively, when inferring each batch of samples. LLM-shearing [50] found that the pruned model has an imbalanced performance across different domains and tasks, and proposed to dynamically load data batch from each domain in proportion to its rate of performance reduction in that domain. K-Prune [38], a retraining-free structured pruning method, used the knowledge loss to measure the importance of heads in transformers, where the knowledge loss is defined as the faithfulness to the soft labels of the original model. FLAP [3] found that certain channels of hidden state features exhibit structured sample stability, and designed a structured pruning metric to identify whether the output feature map is easy to recover when a column of weight matrix is eliminated. Then, FLAP adds an additional bias term to recover the output feature.
@@
Quantization reduces the precision of weights and activations in a neural network by turning the high bit-width numbers into lower bit-width ones. Quantization can directly diminish memory usage and accelerate computations by using fixed-point arithmetic. It enables faster inference on hardware with limited computational resources. Quantization entails quantization-aware training (QAT [34]) and post-training quantization (PTQ [28]). The former emphasizes the combination of quantization and training. The latter quantizes parameters when the network training is completed, which does not require modifications to the network structure but may induce the extra cost of retraining. Therefore, here we mainly discuss PTQ, which is divided into weight quantization and weight&activation quantization.
@@
LLM.int8() [14] initially employs vector-wise quantization alongside distinct normalization constants for every inner product within the matrix multiplication to quantize most of the features. Moreover, LLM.int8() denotes outlier feature dimensions by a 16-bit matrix multiplication, while over 99.9% of values are still multiplied in 8-bit. OPQ [18] is the layer-wise compression method, which performs quantization layer-by-layer and minimizes the error between the pre-activation by the original full-precision and the quantized weight matrix in each layer. SmoothQuant [52] is a training-free, accuracy-preserving, and general-purpose post-training quantization solution. It smoothens the activation outliers by a mathematical scaling transformation before the normal quantization:
@@
where $X$ is the activation and $W$ is the weight matrix. AWQ [33] highlights that not all weights are equally crucial for large language model (LLM) performance. Only 0.1%-1% of weights are significant; bypassing quantization for these salient weights can greatly reduce quantization loss. To identify these salient channels, they examined activation distributions rather than weight distributions. To avoid inefficient mixed-precision implementations, they analyzed weight quantization errors and determined that scaling up salient channels can decrease their relative quantization error. Atom [60] also designs a customized CUDA kernel that utilizes low-bit tensor cores, besides the common quantization operations such as mixed precision and grouped quantization. Atom improves end-to-end throughput (token/s) by up to $7.73$ times compared to the FP16.
@@
Low-Rank Adaptation (LoRA, [27]) posits that the adjustments in weights during model adaptation exhibit a low ”intrinsic rank.” This approach enables the training of certain dense layers in a neural network indirectly by optimizing the gradients into a low-rank decomposition matrix in those layers during adaptation, while keeping the pre-trained weights unchanged. TensorGPT [53] introduced the approach of tensorizing and decomposing each token embedding, rather than treating the entire embedding matrix as a single entity. It was the first method to apply the tensor train decomposition for model compression, achieving a reduction in the number of parameters by a factor of 2.31.
@@ 2.2 Implicit Neural Representation
Implicit Neural Representation (INR) [37] has emerged as a transformative approach in computer graphics, computer vision, and machine learning. This technique leverages neural networks to represent complex shapes and images without explicitly storing the geometry or pixel data. Early work in this area focused on using neural networks to learn mappings from coordinates to values. For instance, the pioneering work by [37] introduced the concept of using multi-layer perceptrons (MLPs) to represent 3D shapes as continuous functions, allowing for high-resolution representations without traditional mesh-based methods. Recently, the implicit neural representation is increasingly used for data compression, such as COIN [16] and NERV++ [22].
@@ 2.3 Hypernet
The concept of hypernetworks was first articulated by [24], who proposed a framework where a neural network (the hypernetwork) generates the weights of another network. This meta-learning approach allows for dynamic adaptation, enabling a single hypernetwork to cater to multiple tasks by producing tailored weights. The hypernetwork architecture is particularly advantageous for scenarios where training multiple models is computationally expensive or infeasible. Hypernet as a meta-learning method has been successfully applied in a plethora of fields including few-shot learning [43] and domain adaptation [49], to name a few.
@@ 3.1 Hyper-Compression
A small human genome can remarkably encode the development of a human brain to the scale of billions of neurons and trillions of synaptic connections [46, 45]. This observation suggests the existence of an implicit mapping from genotype to phenotype, capable of expanding a limited number of genetic instructions into a vast array of biological substances. Inspired by this efficient genetic encoding, the hypernet [9] uses a small network (genotype, called hypernet) to control the design of the target network (phenotype) including architectures and weight distributions.